the length of a rectangle is 5 inches longer than it is wide. If the area is 126 square inches, what are the dimensions of the rectangle?

Answer by addingup(3677)

Area= L*W= 126 (1st equation)
Width= W
Length = W+5 Let's go back to the 1st equation and substitute for L:
(W+5)*W= 126 (2nd equation)
W^2+5W= 126
W^2+5W-126= 0
(W-9)(W+14)= 0
W-9 = 0 or W+14= 0
W= 9 or W= -14 Discard the negative, go back to the 2nd equation and we have:
(W+5)*W= 126
(9+5)*9= 126
14*9 = 126